OpenStudy (anonymous):

find a10 for the geometric sequence An = 3(4)n-1 A: 262,144 B: 786,432 C: 1,048,579 D: 3,145,728 really need help really lost!!! :(

OpenStudy (anonymous):

Are you sure it looks like this? I can't tell if i'm suppose to subtract 1 first or multiply.

OpenStudy (anonymous):

yes you suppose to multiply

OpenStudy (anonymous):

\[a _{n}=a _{1}\times r ^{n-1}\] That's the formula for geometric sequence.

OpenStudy (cp9454):

B

OpenStudy (anonymous):

Ah ok I got it you wrote it wrong. It's suppose to be An=3 x (4)^n-1 Plug and chug 3 x (4)^9= a(10) Do 4^9 first and it equals = 262144 Now multiply that by 3 and it's 786432

OpenStudy (anonymous):

thnkz so much I was lost but now I got the solving down thnkz CP & WORDGEEK :)